August 29, 2016, Analysis Working Group

From GlueXWiki
Jump to: navigation, search

Meeting Time and Place

The meeting will be on Monday August 29, 2016 at 2:00 pm EDT. For those people at Jefferson Lab, the meeting will be in room F224-225.

Meeting Connections

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 115815824.
  3. To join via phone, use one of the following numbers and the Conference ID: 115815824.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. More information on connecting to bluejeans is available.


  1. Announcements
    • Reconstruction Launch
  2. Reconstruction Studies
  3. AOB

Reconstruction Studies


  • Beam Polarization (Justin, Mike, Nathan)
  • Beam Energy (Nathan, Richard)
  • Flux (Justin)


  • CDC Efficiency (Mike)
  • FDC Efficiency (Alex A.)
  • Track Reconstruction (Simon)
  • Tracking Means, Resolution, & Efficiency (Paul, Simon)


  • BCAL
    • Shower Clustering (Will)
    • Neutral Energy & Efficiency (Will)
    • Hadronic Energy & Efficiency (Elton)
  • FCAL
    • Neutral Energy & Efficiency (Jon)
    • Shower Clustering (Matt)
    • Hadronic Energy & Efficiency (Matt)

Particle ID

  • SC Efficiency (Mahmoud)
  • TOF Efficiency (Beni)
  • Particle ID (Paul, All) Run Browser

Channel Reconstruction, Triggering, Acceptance

  • Triggering Efficiency (Alex S., Elton)
  • ρ (Alex A., Justin) [2][3][4]
  • ω→3π (π0 in BCAL, FCAL, both) [5] (Mike S.)
  • φ→2K (Alex B.) [6]
  • Charged 4π (Alex A.) [7]
  • Anti-proton -- Proton photoproduction [8](Reinhard)


  • Reconstruction launch:
    • Improvements to BCAL non-linear corrections coming in soon
    • BCAL Covariance matrix will be ready soon
    • Simon working on FDC track finding improvements
  • Beamline:
    • Beam polarization: Mike Dugger has preliminary TPOL polarization data, will have more soon
    • Flux: Justin is trying to get the tagged flux, working out tagging efficiencies
    • Beam Energy: Fixing the sign discrepancy helped, but there appear to still be more issues. We lined it up manually in the CCDB for the reconstruction launch.
  • Tracking:
    • FDC: Recon efficiency much better now: Psuedo hit efficiencies improved from 70% to 90%, 10% more tracks reconstructed now
    • Preliminary efficiencies show issues in FDC in many-track events, will be discussed at tracking meeting.
  • Calorimetry:
    • BCAL: For high energy Pi0's (e.g. > 6 GeV) the 2 photons are being merged together into a single shower half of the time
    • FCAL: Still working on timing calibration
  • Analysis Channels:
    • Measured resolution is better for 1350 A, but kinfit resolution is about the same for both
    • Yields are higher for all channels (except rho) with 1350 A than 1200 A, cause unknown.
      • The yields for the rho are about the same for both field settings.
    • Will study by trigger type, and will study kinematics, to see where the change in yields might be coming from
    • Simon will look at single-proton + neutral channels to see if magnetic field shows a difference for those channels
    • Reinhard showed that a kinematic fit confidence level cut leaves a clean signal in the ppbar channel.
      • However, there appears to be a phi asymmetry generated by the cut. Justin suggests a looser cut on the kinfit confidence level